# MMBoundary: Advancing MLLM Knowledge Boundary Awareness through Reasoning Step Confidence Calibration

### v1

@@ Abstract
In recent years, multimodal large language models (MLLMs) have made significant progress but continue to face inherent challenges in multimodal reasoning, which requires multi-level (e.g., perception, reasoning) and multi-granular (e.g., multi-step reasoning chain) advanced inferencing. Prior work on estimating model confidence tends to focus on the overall response for training and calibration, but fails to assess confidence in each reasoning step, leading to undesirable hallucination snowballing. In this work, we present MMBoundary, a novel framework that advances the knowledge boundary awareness of MLLMs through reasoning step confidence calibration. To achieve this, we propose to incorporate complementary textual and cross-modal self-rewarding signals to estimate confidence at each step of the MLLM reasoning process. In addition to supervised fine-tuning MLLM on this set of self-rewarded confidence estimation signal for initial confidence expression warm-up, we introduce a reinforcement learning stage with multiple reward functions for further aligning model knowledge and calibrating confidence at each reasoning step, enhancing reasoning chain self-correction. Empirical results show that MMBoundary significantly outperforms existing methods across diverse domain datasets and metrics, achieving an average of 7.5% reduction in multimodal confidence calibration errors and up to 8.3% improvement in task performance [^1] Our code is publicly available at https://github.com/Zhitao-He/MMBoundary..
@@ 1 Introduction
Therefore, in this work, we propose MMBoundary, a reinforced fine-tuning framework for advancing MLLM knowledge boundary awareness by reasoning step confidence calibration. Our method enables the model to express natural language confidence statement for each generated sentence, enhancing reasoning chain self-correction by scaling inference-time. Specifically, we introduce a confidence estimation module that integrates three effective text-based uncertainty methods—namely, length-normalized log probability, mean token entropy, and tokenSAR—and incorporates cross-modal constraint (i.e., CLIPScore) to model the self-rewarded confidence signal from the perspective of its internal states. Then, we propose a mutual mapping between the detected score and predefined confidence statements to achieve two objectives: (1) by inserting confidence statements after the associated knowledge and training the model via supervised learning, we enable the model to naturally generate natural language statements for each sentence, similar to human expression; (2) by integrating internally detected confidence scores and those converted from model expressed statements into the reward modeling for reinforcement learning, we can achieve further confidence calibration, reducing the inaccuracy of model-expressed confidence. Moreover, we annotate the reference reasoning chains of training data to facilitate rigorous evaluation of MLLMs’ knowledge at different reasoning levels, and incorporate model knowledge calibration into the reward modeling, encouraging MLLMs to faithfully express confidence while improving response quality.
@@
  • We propose to integrate both textual and cross-modal self-rewarding signals for confidence estimation. Beyond supervised fine-tuning for initial confidence expression warm-up, we introduce a reinforcement learning stage with multiple reward functions to align model knowledge and calibrate confidence, enhancing self-correction in reasoning chains.
@@ 4.3 Evaluation Metrics
(3) Human Evaluation: Automated model evaluation may not accurately capture the subtle differences between different responses Goyal et al. (2022); Ho et al. (2022); He et al. (2023). Therefore, we conduct additional manual evaluation. We provide a panel of three graduate students with 50 random entries from each setting, asking them to evaluate whether each entry meets the following criteria and to give a score from 1 to 10, consistent with (Xu et al., 2024). 1) Faithful: whether the response faithfully expresses the confidence; 2) Concise: whether the response conveys necessary information clearly and without excess; 3) Granularity: whether the response contains confidence estimates for distinct knowledge. The final result is the average score of these criteria.
@@ 6 Related Work
Hallucinations and Uncertainty Estimation. In VLMs, hallucinations refer to model responses that are misaligned with the visual modality Chen et al. (2024); Liu et al. (2024c), which can arise due to insufficient capabilities in visual perception and knowledge reasoning. Various efforts have been made to evaluate hallucination in VLMs Liu et al. (2024a); Gunjal et al. (2024); Zhang et al. (2024b); Wu et al. (2024); Xu et al. (2025). As a fundamental approach to detecting model hallucination, uncertainty estimation (UE) has long attracted significant attention, falling into two main types: black-box and white-box. Black-box methods only require the generated text, and most of these methods are based on self-consistency Fomicheva et al. (2020); Kuhn et al. (2023); Lin et al. (2023), inspired by its success in textual domain reasoning Wang et al. (2025); He et al. (2024c); Jin et al. (2024). White-box methods rely on access to logits and internal layer outputs. They encompass information-based, density-based and sample diversity-based approaches Malinin and Gales (2020); Kadavath et al. (2022); Vazhentsev et al. (2023); Kuhn et al. (2023); He et al. (2024b); Fadeeva et al. (2024); Duan et al. (2024). Instead of relying on self-consistency prompting, we propose leveraging the model’s internal states to quantify the confidence.
Confidence Calibration of Language Model. Existing research has highlighted the tendency of LLMs to fabricate information when faced with unknown questions Hu et al. (2023); Amayuelas et al. (2023); Liu et al. (2024b); Huang et al. (2025); Fan et al. (2025). As a result, increasing attention has been directed toward enhancing the models’ awareness of their knowledge boundaries and enabling them to express their confidence in outputs when encountering uncertainty Lin et al. (2022); Xiong et al. (2023); Yang et al. (2023); Lyu et al. (2024); Xu et al. (2024); Zhang et al. (2024a). Zhou et al. (2023a) empirically finds that injecting uncertainty expressions into prompts significantly increased the accuracy of GPT-3 responses and improved calibration scores. Zhang et al. (2024a) introduce R-tuning to encourage LLMs to express "certain/not certain". Xu et al. (2024) goes further to teach the model to express more fine-grained confidence estimates along with self-reflective rationales. However, these methods focus solely on the entire response, which can lead to incorrect answers with high confidence. Therefore, we propose MMBoundary to train VLMs to express fine-grained confidence estimates for each reasoning step during inference, enhancing reasoning chain self-correction.
@@ 7 Conclusion
Our framework aims to enable MLLMs to autonomously generate natural language confidence statements during inference, enhancing reasoning chain self-correction. A limitation of this work is that our proposed method involves using the model’s internal states and uncertainty methods to assess the model’s confidence. However, more research is needed to determine whether uncertainty methods can accurately reflect the model’s confidence in its output. Ablation experiments on the uncertainty methods indicate that the four carefully selected methods provide gains for the model. Additionally, we explore the correlation between the proposed internal confidence estimation method and the self-consistency method. The results show that our metric, without requiring multiple samples, achieves performance comparable to methods that rely on multiple samples. To further advance our method, our future work will concentrate on the following areas: Firstly, we aim to incorporate multimodal self-correction mechanisms He et al. (2024a), leveraging diverse data sources to augment the model’s capabilities. Secondly, we plan to explore synthetic data pre-training techniques Qin et al. (2025) to address data scarcity and improve the model’s generalization ability across various project scenarios. Finally, we are committed to expanding the model’s capacity for tool integration and collaboration Qin et al. (2024); Yuan et al. (2024), enabling it to intelligently utilize a broader range of external resources for comprehensive project assistance.
@@ Appendix A The Value-Statement Mapping Table
Table 4: The confidence score-statement mapping table. The five scores correspond to uncertain, slightly uncertain, moderately confident, highly confident, and fully confident. We preset 40 confidence statements for each score.
@@ D.1 The Multi-granular Expected Calibration Error (MECE)
D.2 AUROC
@@
D.3 Reasoning Chian F1 Score
We use the Reasoning Chain F1 score Ho et al. (2022) to evaluate the quality of the reasoning chains generated by the model. We compare the knowledge contained in the predictions and references. First, we split the predicted and reference chains into “steps” by sentence. We then compute a matrix of pairwise similarity scores before using a threshold to classify “matches.” Since a single predicted sentence may contain multiple reference knowledge, we keep separate counts of precise predicted sentences and covered reference sentences. These counts are then micro-averaged to calculate the overall precision, recall, and F1 scores for the test set:
@@ D.3 Reasoning Chian F1 Score
Drawing on the study of (Ho et al., 2022), we select a large RoBERTa model (cross-encoder/stsb-roberta-large) with a similarity threshold of 0.64.

### v2

@@ Abstract
In recent years, multimodal large language models (MLLMs) have made significant progress but continue to face inherent challenges in multimodal reasoning, which requires multi-level (e.g., perception, reasoning) and multi-granular (e.g., multi-step reasoning chain) advanced inferencing. Prior work on estimating model confidence tends to focus on the overall response for training and calibration, but fails to assess confidence in each reasoning step, leading to undesirable hallucination snowballing. In this work, we present MMBoundary, a novel framework that advances the knowledge boundary awareness of MLLMs through reasoning step confidence calibration. To achieve this, we propose to incorporate complementary textual and cross-modal self-rewarding signals to estimate confidence at each step of the MLLM reasoning process. In addition to supervised fine-tuning MLLM on this set of self-rewarding confidence estimation signal for initial confidence expression warm-up, we introduce a reinforcement learning stage with multiple reward functions for further aligning model knowledge and calibrating confidence at each reasoning step, enhancing reasoning chain self-correction. Empirical results show that MMBoundary significantly outperforms existing methods across diverse domain datasets and metrics, achieving an average of 7.5% reduction in multimodal confidence calibration errors and up to 8.3% improvement in task performance [^1] Our code is publicly available at https://github.com/Zhitao-He/MMBoundary..
@@ 1 Introduction
Therefore, in this work, we propose MMBoundary, a reinforced fine-tuning framework for advancing MLLM knowledge boundary awareness by reasoning step confidence calibration. Our method enables the model to express natural language confidence statement for each generated sentence, enhancing reasoning chain self-correction by scaling inference-time. Specifically, we introduce a confidence estimation module that integrates three effective text-based uncertainty methods—namely, length-normalized log probability, mean token entropy, and tokenSAR—and incorporates cross-modal constraint (i.e., CLIPScore) to model the self-rewarding confidence signal from the perspective of its internal states. Then, we propose a mutual mapping between the detected score and predefined confidence statements to achieve two objectives: (1) by inserting confidence statements after the associated knowledge and training the model via supervised learning, we enable the model to naturally generate natural language statements for each sentence, similar to human expression; (2) by integrating internally detected confidence scores and those converted from model expressed statements into the reward modeling for reinforcement learning, we can achieve further confidence calibration, reducing the inaccuracy of model-expressed confidence. Moreover, we annotate the reference reasoning chain of training data to facilitate rigorous evaluation of MLLMs’ knowledge at different reasoning levels, and incorporate model knowledge calibration into the reward modeling, encouraging MLLMs to faithfully express confidence while improving response quality.
@@
  • We propose to integrate both textual and cross-modal self-rewarding signals for confidence estimation. Beyond supervised fine-tuning for initial confidence expression warm-up, we introduce a reinforcement learning stage with multiple reward functions to align model knowledge and calibrate confidence, enhancing self-correction in reasoning chain.
@@ 4.3 Evaluation Metrics
(3) Human Evaluation: Automated model evaluation may not accurately capture the subtle differences between different responses Goyal et al. (2022); Ho et al. (2022); He et al. (2023). Therefore, we conduct additional manual evaluation. We provide a panel of three graduate students with 50 random entries from each setting, asking them to evaluate whether each entry meets the following criteria and to give a score from 1 to 10, consistent with (Xu et al., 2024). 1) Faithful: whether the response faithfully expresses the confidence; 2) Concise: whether the response conveys necessary information clearly and without excess; 3) Granular: whether the response contains confidence estimates for distinct knowledge. The final result is the average score of these criteria.
@@ 6 Related Work
Hallucinations and Uncertainty Estimation. In MLLMs, hallucinations refer to model responses that are misaligned with the visual modality Chen et al. (2024); Liu et al. (2024c); huang2025adactrladaptivecontrollablereasoning, which can arise due to insufficient capabilities in visual perception and knowledge reasoning. Various efforts have been made to evaluate hallucination in MLLMs Liu et al. (2024a); Gunjal et al. (2024); Zhang et al. (2024b); Wu et al. (2024); Xu et al. (2025). As a fundamental approach to detecting model hallucination, uncertainty estimation (UE) has long attracted significant attention, falling into two main types: black-box and white-box. Black-box methods only require the generated text, and most of these methods are based on self-consistency Fomicheva et al. (2020); Kuhn et al. (2023); Lin et al. (2023), inspired by its success in textual domain reasoning Wang et al. (2025); He et al. (2024c); Jin et al. (2024). White-box methods rely on access to logits and internal layer outputs. They encompass information-based, density-based and sample diversity-based approaches Malinin and Gales (2020); Kadavath et al. (2022); Vazhentsev et al. (2023); Kuhn et al. (2023); Fadeeva et al. (2024); Duan et al. (2024). Instead of simply relying on self-consistency prompting, we leverage the internal state of model to quantify the confidence of each reasoning step.
Confidence Calibration of Language Model. Existing research has highlighted the tendency of LLMs to fabricate information when faced with unknown questions Hu et al. (2023); Amayuelas et al. (2023); Liu et al. (2024b); Huang et al. (2025); Fan et al. (2025). As a result, increasing attention has been directed toward enhancing the models’ awareness of their knowledge boundaries and enabling them to express their confidence in outputs when encountering uncertainty Lin et al. (2022); Xiong et al. (2023); Yang et al. (2023); Lyu et al. (2024); Xu et al. (2024); Zhang et al. (2024a). Zhou et al. (2023a) empirically finds that injecting uncertainty expressions into prompts significantly increased the accuracy of GPT-3 responses and improved calibration scores. Zhang et al. (2024a) introduce R-tuning to encourage LLMs to express "certain/not certain". Xu et al. (2024) goes further to teach the model to express more fine-grained confidence estimates along with self-reflective rationales. However, these methods focus solely on the entire response, which can lead to incorrect answers with high confidence. Therefore, we propose MMBoundary to train MLLMs to express fine-grained confidence estimates for each reasoning step, enhancing reasoning chain self-correction.
@@ 7 Conclusion
Our framework aims to enable MLLMs to autonomously generate natural language confidence statements during inference, enhancing reasoning chain self-correction. A limitation of this work is that our proposed method involves using the model’s internal states and uncertainty methods to assess the model’s confidence. However, more research is needed to determine whether uncertainty methods can accurately reflect the model’s confidence in its output. Ablation experiments on the uncertainty methods indicate that the four carefully selected methods provide gains for the model. Additionally, we explore the correlation between the proposed internal confidence estimation method and the self-consistency method. The results show that our metric, without requiring multiple samples, achieves performance comparable to methods that rely on multiple samples. To further advance our method, our future work will concentrate on the following areas: Firstly, we aim to incorporate multimodal self-correction mechanisms He et al. (2024a), leveraging diverse data sources to augment the model’s capabilities. Secondly, we plan to explore synthetic data pre-training techniques Qin et al. (2025) to address data scarcity and improve the model’s generalization ability across various project scenarios.
@@ Appendix A The Value-Statement Mapping Table
Table 4: The confidence score-statement mapping table. The five scores correspond to uncertain, slightly uncertain, moderately confident, highly confident, and fully confident. We preset 40 statements for each score.
@@ D.1 The Multi-granular Expected Calibration Error (MECE)
D.2 Area Under the Receiver Operating Characteristic curve (AUROC)
@@
D.3 Reasoning Chain F1 Score
We use the Reasoning Chain F1 score Ho et al. (2022) to evaluate the quality of the reasoning chain generated by the model. We compare the knowledge contained in the predictions and references. First, we split the predicted and reference chains into “steps” by sentence. We then compute a matrix of pairwise similarity scores before using a threshold to classify “matches.” Since a single predicted sentence may contain multiple reference knowledge, we keep separate counts of precise predicted sentences and covered reference sentences. These counts are then micro-averaged to calculate the overall precision, recall, and F1 scores for the test set:
@@ D.3 Reasoning Chain F1 Score
Drawing on the study of Ho et al. (2022), we select a large RoBERTa model (cross-encoder/stsb-roberta-large) with a similarity threshold of 0.64.